You have already invested $550 in a stock with an annual return of 11%. How much of an additional $1,100 should be invested at 12% and how much at 6% so that the total return on the entire $1,650 is 9%? (Round each answer to the nearest cent.)
Answer by richwmiller(17219)

I just did one that is almost identical.
http://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.911728.html
z+x+y=1650
.11*550+0.06*x+0.012*y=.09*1650
60.50+0.06*x+0.012*y=148.50
1650-550=1100
148.50-60.50=88
Now we have two equations in two unknowns
x+y=1100
0.06*x+0.012*y=88